Sales Outcome
Mean 0.481
Standard Error 0.015808
Median 0
Mode 0
Standard Deviation 0.499889
Sample Variance 0.249889
Range 1
Minimum 0
Maximum 1
Sum 481
Count 1000

For all the tests listed below assume that the significance value a=0.05, unless otherwise stated.

5. Perkins,CMO informed the board that they win at least 50% of the sales leads that they receive. Use an appropriate hypothesis testing procedure to check whether the proportion of leads that they receive. Use an appropriate hypothesis testing procedure to check whether the proportion of leads won by WSES is more than 50%.

6. Jackson, who works the product line "learns", claims that the probability of winning a sales lead for the product "learns" is more than that of "Fins". Is there statistically significant evidence in favor of Jacksons claim?

7. John also claims that the average sales value of "Learns" projects is higher than that of "Fins" projects. Check whether John is correct at 5% significance.

8. Jack believed that the sales conversions are different for different products as well as different geographical locations. Check the validity of Jack's belief using an appropriate hypothesis test to check the validity of this claim by making the following 3 groups:
Answer #1

5. The hypothesis being tested is:

H0: p = 0.50

Ha: p > 0.50

= 0.481

n = 1000

The test statistic, z = (p̂ - p)/√p(1-p)/n = (0.481 - 0.50)/√0.50(1-0.50)/1000 = -1.20

The p-value for z = -1.20 is 0.8853.

Since the p-value (0.8853) is greater than the significance level (0.05), we cannot reject the null hypothesis.

Therefore, we cannot conclude that the proportion of leads won by WSES is more than 50%.